A body is weighed by a spring balance to be 1.000 kg at the north pole. How much will it weigh at the equator? Account for the earth’s rotation only.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. Let g’ be the acceleration due to gravity at equation & that of pole = g g’ = g ω^2 R = 9.81 – (7.3 * 10^-5)^2 * 6400 * 10^3 = 9.81 – 0.034 = 9.776 m/s^2 Mg’ = 1 kg * 9.776 m/s^2 = 9.776 N or 0.997 kg The body will weight 0.997 kg at equator